OpenStudy (anonymous):

factor the binomial of y^15-8z^9

OpenStudy (anonymous):

(y^5-2 z^3) (y^10+2 y^5 z^3+4 z^6)

OpenStudy (mathstudent55):

This is a difference of cubes. y^15 = (y^5)^3, so it's the cube of y^5 8z^9 = (2z^3)^3, so it's the cube of 2z^3 Use the factoring of a difference of cubes.